OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (2,12) (6,11)

OpenStudy (texaschic101):

are you familiar with the slope formula ?

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) <== slope formula (2,12)....x1 = 2 and y1 = 12 (6,11)...x2 = 6 and y2 = 11 now we sub slope = (11 - 12) / (6 - 2) slope = -1/4 understand ?

OpenStudy (anonymous):

Not completely but more than i did before.